An immersion heater has a resistance of 20 ohms and is used with 240 V mains supply. What is its power output?
1
Expert's answer
2020-09-17T14:28:51-0400

Let's find the current in immersion heater. According to Ohm's law:

"\displaystyle V = IR \; \; \Rightarrow \; \; I = \frac{V}{R}"

The power output is

"\displaystyle P = VI = \frac{V^2}{R} = \frac{240^2}{20}= 28\;800 \; \textrm{W} = 28.8 \; \textrm{kW}"

Answer: "28.8" kW